### v1

@@ 6 Conclusion and future work
We presented a computationally efficient RL framework for MPC policies that enables computationally low-cost second-order policy optimization. The key contribution is a Gauss-Newton approximation of the deterministic policy Hessian that avoids the computation of second-order policy derivatives while still enabling superlinear convergence. The resulting approximation is well-suited for MPC policies as it reduces the amount of expensive agent-environment interactions drastically due to its improved convergence behavior, while it simultaneously avoids to compute the computationally-expensive second-order NLP sensitivities. Furthermore, we proposed a momentum-based exponentially moving averaging scheme for Hessian estimation that improves robustness against noise in local curvature samples. The proposed method improves stability during training and enables faster updates, which is beneficial for sample efficiency.
We rigorously proved the superlinear convergence behavior of the proposed Gauss-Newton approximation. This was confirmed empirically on an analytical example and a nonlinear CSTR benchmark. Results show that the Gauss-Newton approximation in combination with momentum-based averaging significantly reduces the number of parameter updates compared to state-of-the-art first-order methods. In addition, the proposed approach also reduces the total wall-clock time for training compared to first-order and approximate policy Hessian methods while still maintaining robust learning behavior.

### v2

@@ 6 Conclusion and future work
We presented a computationally efficient RL framework for MPC policies that enables computationally low-cost second-order policy optimization. The key contribution is a Gauss-Newton approximation of the deterministic policy Hessian that avoids the computation of second-order policy derivatives while still enabling superlinear convergence. The resulting approximation is well-suited for MPC policies as it can reduce the amount of expensive agent-environment interactions drastically due to its improved convergence behavior, while it simultaneously avoids to compute the computationally-expensive second-order NLP sensitivities. Furthermore, we proposed a momentum-based exponentially moving averaging scheme for Hessian estimation that improves robustness against noise in local curvature samples. We embed the Gauss-Newton Hessian approximation and the momentum-based averaging scheme in a trust-region constrained optimization problem. The proposed method improves stability during training and enables faster updates, which is beneficial for sample efficiency.
We rigorously proved the superlinear convergence behavior of the proposed Gauss-Newton approximation. This was confirmed empirically on an analytical example and a nonlinear CSTR benchmark. Results show that the Gauss-Newton approximation can reduce the number of parameter updates compared to state-of-the-art first-order methods and state-of-the-art deep RL approaches. In addition, the proposed approach also reduces the total wall-clock time for training for increasingly more complex parameterized policies compared to more elaborate approximate policy Hessian methods while still maintaining robust learning behavior. The proposed Gauss-Newton approach also shows resilience with respect to prescaling of the policy parameters.